UT 08 Sep 2021 - Night #7 of Restart




Night Staff:

OSA: D.G. Huerta

ISp: E. Solheid

ISA: O. Kuhn

Technical Wizard: J. Hill (Tucson)

Mtn Manager: M. Wagner

Sunset: 18:41

Twlight Ends: 20:00

Twilght Starts: 04:35

Sunrise: 05:54

Starting Config: LUCI-1 pending Eng. Work

Night Log:

UT 01:29 - Engineering work done to address mask issue in MODS-2 (now corrected), and replace the LBC-R controller with Controller #4 (in progress). Likely opening will be delayed.

MODS checkout continued...

Longslit acquisition test

02:55 We're open and collimating near the first target. It is at low airmass, 40 deg elevation. Seeing here is ~1.2".

03:01 acqBinoMODS PDS456_pa55_acqblue.txt PDS456_pa55_acq.txt

On MODS1: modsAlign -y 11 210903M8.01j.fits 210903M8.01k.fits --> mods1b 0001, pretty well-centered

On MODS2: modsAlign -y 9 210903M8.07q.fits 210903M8.07r.fits --> mods2r 0001, pretty well-centered though a tad to the right. Not tweaking the centering since clouds are coming.

03:12 execBinoMODS PDS456_pa55_obs.txt to take a set of 300-sec dual grating exposures.

Dither Test

03:47 Dither test (pre-transform). This test dithers in +/- Y with the pointing mask in place to test the alignment of the dithers with the mask.

Test also sending offsets from the GUI because we need to adjust the positions of the stars so they are in the pointing mask at both dither positions 30" apart. Based on m1r: 004 and m2r 0004 I offset by 2.5, -2.5 on MODS1 and 7.4 -3.? on MODS2

04:23 25 and 24 - guide star was off of hotspot and m2r 25 is trailed, m2b 24 is less so but similar.

Files m1b: 5-34 and m1r: 5-33. m2b: 5-33 m2r: 5-34 (double check numbers)

Blind Offset test

04:44 acqBinoMODS Gaia20p25_BlindOffset_blue.txt Gaia20p25_BlindOffset.txt

Center on star at 19:59:57.729 25:58:25.59 and offset by 100.37 168.310 to star at 20:00:05.112 25:01:13.90

04:46 Closing for threatening clouds. With MODS, the blind offset and non-sidereal tests remain.

Alessandro and Yang adjusted the bias voltages to get bias levels of 300 ADU on each chip of LBCR which is now using controller #4.

LBC Gain/Linearity data Closed-Dome

Reconfiguring to the LBCs to collect data for gain/linearity measurements.

05:55 Starting data collection for linearity/gain using the script on robs in ~/kuhn/23May2021_Linearity.xml. It takes a series of images in V-Bessel on Blue and i-SLOAN on Red with the Chamber footlights & left+right stairwell lights.

06:40 Conditions improving. Preparing to open again. Since we have LBCs in the beam, we'll run a science-like OB before reconfiguring.

LBC collimation and quick B+R guided+dithered observation (B+R)

07:10 Open. Slewing to MCG+06 copointing.

07:17 dohybrid - there was a delay in images appearing in /lbt/data/new/ and dohybrid paused - eventually they appeared and I entered <CR> to continue the analysis. Converged in 2 FPIA iterations

07:22 Start OB with guiding + dithering at B+R.

MODS tests part 2

07:32 Reconfiguring to MODS

07:53 Checking pointing and collimation. Object is low and DX has some strong coma & astigmatism.

MODS - Blind Offset Test

07:53 acqBinoMODS Gaia20p25_BlindOffset_blue.txt Gaia20p25_BlindOffset.txt

no star found on left side, hit "r" to retry worked. A glitch in GCS?

MODS2 Red IMCS failed to lock, retry worked. Elevation = 38 deg, RDG rotator = 273 deg.

m1b: 35 & 36 --> 35 (slit image) is messed up. star+slit is imaged multiple times

m2r: 35 & 36 --> 35 slit image is also messed up on MODS2 - looks like a bias.

Repeating the instrument config on both sides via acqMODS --modsN -i and then -a to take the images.

m1b: 37 & 38 -->

m2r: 37 & 38 -->

lost guide star - clouds... lost the guide star during the acquisition. Was the offset completed or not?

Repeat m1b acq images: m1b 39 & 40 --> MASKPOS = EMPTY in 39 where it should be STOW. Edited header in copy of file.

m2r: 37 & 38 --> offsetxy 3.755 9.290 rel (offset made, but guide star fading) --> 39 (lost guide star) and 41.

m1b: 39 & 40 --> offset had been made so star was spot on and made only a very small adjustment. --> 41 (losing guide star and mask is not in FP)

With clouds and the need to repeat, I've lost track of the images, but confirmatory images of reference star, m1b 42 and

Confirmatory images look good enough, resuming script --> m1b 43 (and m2r 40 look good).

Send blind offset but copointing limit. Argh! - After the initial errors with the IMCS (MODS1) and bias-like image (MODS2), I sent the scripts monocularly without the --bino flag. Without it, syncoffset is not interpreted.

Non-sidereal imaging with dithers

execBinoMODS 531_Zerlina_8UT.txt

The script takes 3 60-sec exposures at each of 3 dither positions.

WFS BLOCKED: try to shorten exposure time. Rates are not excessive: -15"/hr in ra and -45"/hr in dec

The same script was sent on both sides, but for MODS2, only 1 image at each dither was taken. ?? Stale state flags?

Blind Offset test

Error - copointing limit when I slewed - telescope does not point to fields 8 hr apart!?

09:17 acqBinoMODS Gaia04p25_BlindOffset_blueacq.txt Gaia04p25_BlindOffset.txt

09:32 again - this time having edited blueacq to use the blue camera.

m1b: 53 & 54 --> 55 well-centered

m2r: 47 & 48 -->

09:42 Error - AZ power amplifier failed... presets have been canceled. But in recovery problem arose in EL brake. A TPLC problem. David left a message with Dan Cox.

10:28 Problem resolved - back to blind offset. But robs has slowed down (have MODS, LBC, LUCI and OT all running, though I did before the issue too). ds9 is taking 99% CPU (owned by telescope). But killing that process did not help.

top - 10:28:49 up 21 days, 13:15, 4 users, load average: 1.60, 1.95, 2.05
Tasks: 473 total, 4 running, 469 sleeping, 0 stopped, 0 zombie
%Cpu(s): 7.4 us, 9.9 sy, 0.0 ni, 82.3 id, 0.0 wa, 0.3 hi, 0.1 si, 0.0 st
MiB Mem : 64324.2 total, 477.7 free, 6007.4 used, 57839.0 buff/cache
MiB Swap: 0.0 total, 0.0 free, 0.0 used. 57406.8 avail Mem

PID USER PR NI VIRT RES SHR S %CPU %MEM TIME+ COMMAND
20426 telesco+ 20 0 423800 88408 20540 R 98.7 0.1 7079:40 ds9
7477 root 20 0 11628 6032 4580 R 1.0 0.0 0:00.03 x2gogetstatus
2372 lbto 20 0 21.4g 821780 26368 S 0.7 1.2 1:46.34 java
7000 lbto 20 0 17792 4748 3700 R 0.7 0.0 0:00.73 top
14367 lbto 20 0 867856 491952 86264 S 0.7 0.7 59:54.30 x2goagent
1 root 20 0 172500 15768 9720 S 0.3 0.0 1:22.19 systemd
836 root 20 0 57240 20228 18852 S 0.3 0.0 27:40.95 sssd_nss
2219 lbto 20 0 126004 89168 23960 S 0.3 0.1 0:34.99 ds9
2433 lbto 20 0 19.6g 379952 23284 S 0.3 0.6 0:17.01 java
2467 lbto 20 0 19.6g 382056 23304 S 0.3 0.6 0:15.35 java


10:34 Sending preset from obs2 - robs is still hung. I killed and resumed my x2go session on robs - now it responds. Interestingly (related?), both scripts timed out in setting OBSMODE (I left the GUIs on robs while I killed/resumed my x2go onto it).

race condition behavior - Go - exposure already in progress - on both sides. Stopped. Bounced GUIs.

10:41 Sending again. Aligning. On MODS2 offset from modsAlign ERROR: Command 'offsetxy 3.511 8.314 rel' timed out after 60 seconds. Clicked button on gui but it doesn't seem to be offsetting. Sync preset waiting.

tcs process on MODS2 died. Don't know how it happened. Restarting.

11:05 Trying again. modsAlign done and offsets made successfully this time.

Confirmatory thru-slit images of "ref" star are: M1B 62 and M2R 54. Star looks well centered.

Sync offset -154.780 -202.680 made. Taking thru-slit images of the "target" star: M1B 63 and M2R 54. The target star is bang-on in MODS1 but a bit on the left and up on MODS2.

LUCI1 tests

11:22 Reconfiguring to LUCI1

Long slit with G210

11:40 After pointing & collimation, sent preset to NGC 455 for a long slit test

Acq images are 16&17, 18 is the mask

Sent offset -0.2309" 0.1237" to center galaxy in slit. 19 is the confirmatory throught-slit. It looks centered.

Starting the spectroscopic exposures using G210 @ 2.3 mic and 0.5" slit: 0020 -> 0022 (set pause to stop early as it's getting light)

Repeating target observations, but this time at G200 @1.93 mic HK (100 sec): 0023-0024

Now at G200 zJ @1.17 mic (250sec): 0025-0026 We're well into twilight.

12:22 Background seen in WFS

12:23 Telluric for NGC 455 --- let's give it a try even though it looks very bright and I don't see any stars on the allsky. The guide star is found, and the WFS images look ok. Seeing 0.65"

The star, BD+18 338A is a double (simbad) but the orientation of the two stars in the simbad chart is different from what I see - it is flipped in X. Our PA=0. Not sure I chose the correct star - I chose the one at top right. Puzzling - a ghost image through crossed filters, but both images seem about equally bright.

12:32 Observing it at G210 2.30 mic: 0030, 0031 and 0032. guiding is lost on image 0032 but it was for the first two.

12:38 Configuring for imaging to check HeI +J image - was the star double or a ghost? - by making one or two offsets.

0033 (0,0)

0034 (20,20 detxy offset) we were still in active mode and the offset failed. David turned off guiding and I repeated the offset

0035 (-50,-40)

It's a double star (later in the morning, I saw the light --- the double is not resolved by Simbad and there 2nd star in the 4.5' field of the Simbad image that is to the SW of BD+18. We see that star, too, in the LUCI images - it's about 1 arcmin away.

12:43 Skyflats at blank field 02+13.

Ks are bright - 36k counts ... 0041

H 0042 --- wrong way for morning

BrG 0043-0047 --- 6000 cnts

FeII 0048-0052 32000 in center to 29K around edges

H2 0053-0057 20000-23000 counts

BrG 0058-0062 21000-26000 counts

13:02 David is closing up.
  • Putting blind+blind filters in LUCI1, no mask, N3.75 camera.
  • Putting MODS to sleep. I will wait for the headers to be fixed before taking cals.
  • Taking biases with the LBCs, starting at 13:11

.

Proposed Plan of Action:

  • LBC
    • Photometric standards (clear any seeing)
    • Spectro-photometric standards (clear any seeing)
    • Twilgiht Flats
    • TMS w/LBC
      • Active
      • Active + thermal corrections
    • Pointing Model (poor weather) - higher priority to address rotational trailing
      • can be done if controllers are swapped on LBC-R
      • absent LUCI-2 which could affect results
    • Check LBCs in parallel with other instruments
      • take images and visually check
  • MODS
    • Re-take transform data with MODS-1 Blue to address issues with MODS-1 Red header corruption
    • Test Science OB
      • longslit
      • Non-Sidereal (longslit or imaging)
    • Skyflats
    • IT 8315 GCS and IDL have opposite field correction - wait for feedback from J. Hill and D.Miller - low for now
    • Blind Offset Test - high
    • Pointing Model (poor seeing and clouds)
      • with AGWs - if either MODS not back to full status
      • absent LUCI-2 which could affect results
  • LUCI
    • Longslit G200 check (sidereal w/dithers) - high
    • Longslit G210 check (sidereal w/dithers) - high
    • NS longslit (more to check dithering stays in slit OK than checking grating)
    • Imaging dither in-field (low, big dithers worked OK)
    • Twilight Flats
    • Poniting Model (thick clouds, poor seeing)
      • with AGWs
      • absent LUCI-2 which could affect results
    • Photometric stars (clear any seeing)


Current Status

  • LBCs
    • LBC-Blue - available
    • LBC-Red - available
  • MODS
    • MODS-1 - issues with M1RC headers & filenumbers short 1 digit
      • Imaging OK both sides, spectroscopy blue only for acquisition
      • AGW-5 - available
    • MODS-2 - available
      • AGW-6 - available
  • LUCI-1 - available
    • AGW-1 available
  • LUCI-2 offline
    • AGW-2 available
  • PEPSI - offline Software environmental system issues - earliest Monday
    • AGW-7/8 available


Closed Dome Activities

  • MODS-1 Full set of Cals
    • Biases
      • 1x1
      • 1x2
      • 2x2
    • Grating
      • Dual
      • Red
      • Blue
    • Imaging
      • Dual
      • Red
      • Blue
    • Prism
      • Dual
      • Red
      • Blue
  • MODS-2 Cals
    • Bias
      • 1x2
      • 2x2
    • Prism
      • Dual
      • Red
      • Blue
  • LUCI-1 Cals
    • Imaging N375 Flats
    • Narrow, medium-band filters
    • Broad zJHKKs and HeI completed
  • ARGOS Cal-Unit test for M3 alignment and pupil wobble checkout (alternate to on-sky method)
  • LUCI-1 Only Calibrations
    • Flexure data for G150
      • /home/dthomspons/Luci/FlexureCompensation
      • README - Calib15 - 2hrs
      • Ks filter, Argon lamp, optical sieve mask
      • take dark frame
      • Manual Control MCSGUI for elevations
    • Wavelength data for G150
      • G150 Ks, Argon, take image through slit, then put in grating, line should show up where the slit is located (within a few pixels of the image of the slit)
      • Check a few lines (2 or 3)


-- BarryRothberg - 07 Sep 2021
Topic revision: r10 - 09 Sep 2021, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback